What is the formal charges on atoms: PF3? What is the hybridization ion on central atom?
Answer #1

PF3 FC on P = 5-3-2 = 0 FC on each F = 7.1-6 = 0 hybridization of P = sp3 4 electronic groups (3 bonding + 1 nonbonding) pres